# Unveil Inversion and Invariance in Flow Transformer for Versatile Image Editing

### v3

@@ Abstract
Leveraging the large generative prior of the flow transformer for tuning-free image editing requires authentic inversion to project the image into the model’s domain and a flexible invariance control mechanism to preserve non-target contents. However, the prevailing diffusion inversion performs deficiently in flow-based models, and the invariance control cannot reconcile diverse rigid and non-rigid editing tasks. To address these, we systematically analyze the inversion and invariance control based on the flow transformer. Specifically, we unveil that the Euler inversion shares a similar structure to DDIM yet is more susceptible to the approximation error. Thus, we propose a two-stage inversion to first refine the velocity estimation and then compensate for the leftover error, which pivots closely to the model prior and benefits editing. Meanwhile, we propose the invariance control that manipulates the text features within the adaptive layer normalization, connecting the changes in the text prompt to image semantics. This mechanism can simultaneously preserve the non-target contents while allowing rigid and non-rigid manipulation, enabling a wide range of editing types such as visual text, quantity, facial expression, etc. Experiments on versatile scenarios validate that our framework achieves flexible and accurate editing, unlocking the potential of the flow transformer for versatile image editing.
@@ 1 Introduction
To address the aforementioned issues, we systematically investigate flow inversion with the Euler method and invariance modeling in MM-DiT. For inversion, we reveal that the reverse ordinary differential equation (ODE) with the Euler method has a similar structure to DDIM inversion but is more affected by the approximation error between two consecutive states. This error makes Euler inversion diverge more from the original image than DDIM inversion in diffusion and significantly deteriorates the editing ability even if the image can be recovered by optimization. So, we proposed a two-stage flow inversion. First, we use the fixed-iteration technique to reduce the approximation error when estimating the flow velocity and get a basic inversion with the Euler method. This makes inversion approximates the authentic generation process. Then based on such inversion, we only need to add mild compensation at each denoising step to recover the exact original image. Since the compensation at each denoising step is small, the inversion trajectory is close to the model’s original domain, thereby preserving the editing ability.
@@ 2 Related Work
Image editing via diffusion models. Leveraging the large prior of the text-to-image diffusion model, earlier editing methods fine-tune the model with the original image and the text prompt [17]. Later approaches mainly fall into three categories. Truncation methods analyze latent noise predictions of the model [1, 18] and filter out undesired editing parts to preserve the non-target regions. The instruction-training methods [3, 42, 7, 31, 23, 9] require collecting a triplet of original and edited images and the editing instruction to re-train the diffusion model to be an editing model, which can be difficult to scale up. The inversion-based methods [40, 10, 26, 35, 4] are tuning-free and first convert the original image into the initial noisy latent and regenerate the edited image with the edited prompt while preserving the invariance (i.e., unchanged content from the original image). Others further improve the applicable ability by combining the large language model [14, 19] and enrich the editing forms [32, 41, 20, 22].  Hu et al. [13] research the latent space for flow-based image editing but the backbone is U-ViT yet the large-scale DiT model, and does not discuss the flow inversion.
@@
Inversion of diffusion for real image editing. The inversion that converts the image into the initial noisy latent is pivotal in tuning-free editing. One branch leverages DDIM [33] to invert the ODE process. However, the classifier-free guidance (CFG) in editing deteriorates the accumulated error in reversing ODE. Various mitigation solutions explore optimizing the null-text embedding [25], adding corrections [16], better estimation of latents [6], and negative prompt [24]. Another branch adopts the stochastic differential equation (SDE) instead of ODE to mitigate the error by manipulating random noise [2, 15]. Besides these heuristic solutions, others explore the mathematical exact inversion by reformulating the sampling process [36, 12, 37]. Besides these nonparametric solutions, TurboEdit [38] learns the inversion process with a neural network to reduce the inversion time. Unlike these diffusion-oriented inversions, we specifically unveil the deficiency of flow inversion and study efficient inversion based on the velocity field of flow models.
@@
Invariance control in diffusion-based image editing. The invariance control preserves the original image’s unedited contents during editing. The instruction-based methods explicitly add the original image features as the condition and retrain the T2I model into a text-guided image-to-image (TI2I) model [3, 7]. In the tuning-free paradigm, attention-based methods [10, 35] inject attention maps from the original image to preserve the original contents but may struggle with non-rigid editing such as layout and pose. Later methods [4, 39] propose the mutual-self attention that injects K and V values in the later diffusion process to adapt to the layout change but this deteriorates the fidelity for large changes. The refinement approach [1, 18] filters out components of the predicted noise corresponding to the non-target regions to preserve the non-target regions. However, these methods require careful tuning filter for different editing types. Moreover, these approaches are mainly based on diffusion and U-Net. In contrast, to take full advantage of the generation ability of the flow transformer, we systematically investigate inversion for rectified flow and flexible invariance control based on MM-DiT to support rigid and non-rigid editing, enabling versatile editing types.
@@
Figure 2: Framework of rectified flow inversion and editing. Left: The two-stage inversion for rectified flow. A basic inversion trajectory is first constructed to resemble the generation process, and then mild compensation is added to velocity to recover the image exactly. Right: Invariance control with text feature replacement within AdaLN during sampling. In the target branch, the unchanged text features are replaced with features from the original image while the edited text features remain intact.
@@ 3 Preliminary
We aim to design a tuning-free framework based on the flow transformer for versatile image editing. In the following, we first briefly introduce the rectified flow and DDIM inversion in diffusion in Sec 3. Then, we analyze the deficiency of Euler inversion in rectified flow and propose the two-stage flow inversion in Sec 4. In Sec 5, we present the image invariance investigation in MM-DiT centered on text features in AdaLN. Last, we summarize the full algorithm.
@@
DDIM. DDIM [33] is the implicit probabilistic model trained with DDPM [11] objective and has the corresponding ODE formulation below where $\mathbf{x}_{t}$ is also the noisy data sample, $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$ is the noise prediction, $\mathbf{f}_{\theta}(\mathbf{x}_{t},t)$ is the predicted original data sample.
@@ 4.1 Deficiency and Relation of Euler to DDIM
Essentially, if we rescale $\mathbf{x}_{t}$ with $\frac{\mathbf{x}_{t}}{\sqrt{\alpha_{t}}}$ and rewrite Eq. 5, DDIM can also be treated as the first-order ODE solver as Euler above, and there has been extensive research on optimizing the DDIM inversion in DM. However, despite the similar formulation of Euler and DDIM inversion, the performance of Euler inversion is significantly worse than the DDIM even if the Euler inversion is evaluated on the more powerful Stable Diffusion 3.5 vs SD 1.5 as shown in Figure 3. Such phenomena are not expected and also make optimization methods [25, 16, 2] in diffusion invalid for flow inversion.
@@
We argue this is due to the fact that the Euler inversion for rectified flow transformer suffers more from the approximation error. To reduce it, we need to mitigate the gap between $v_{\theta}(\mathbf{x}_{t+1},t)$ and $v_{\theta}(\mathbf{x}_{t},t)$. Idealy, we expect to use the exact inversion in Eq. 7 to get the initial noisy latent $\mathbf{x}_{0}$ and the transition trajectory $\{\mathbf{x}_{1},...,\mathbf{x}_{t}\}$. Such an inversion trajectory is close to the trajectory in the generating process and requires minimal optimization. Thus, it fits the prior distribution of the model and has good editing ability.
@@ 4.2 Two-stage Flow Inversion
Our principle is to find an initial latent and inversion trajectory that is as close to the authentic generation process as possible, trying to avoid only overfitting the original image. This has two-fold benefits: 1. Editing friendly. The latent and trajectory fit into the prior distribution of the model and text-to-image alignment is authentic so that the image can be easily manipulated by changing the text prompt. 2. Easy to preserve invariance and fidelity. If the inversion trajectory deviates from the model prior, it requires more effort such as self-attention injection to preserve the invariance and fidelity. However, injecting too much self-attention from the original image can hinder the editing. Thus, we propose a two-stage flow inversion. First, we construct a pivotal inversion close to the generating process. Second, we manually add compensation at every timestep to eliminate the left mild error.
@@
Besides, ideally, the velocity $v_{\theta}$ is $\mathbf{x}_{1}-\mathbf{x}_{0}$ according to Eq. 3 that is a stable value. This also strengthens the motivation of fixed-point iteration, and we can use this property to average the velocities to get a more stable velocity prediction. Practically, this is also equivalent to averaging the latent series $\{\mathbf{x}_{t}^{i}\}_{i=1}^{I}$.
@@
1:  Input: An image $\mathbf{x}_{1}$, number of iteration steps $I$, number of inversion steps $T$, velocity model $v_{\theta}$.
@@
6:        $\mathbf{x}_{t-1}^{i}=\mathbf{x}_{t}+(\sigma_{t-1}-\sigma_{t})v_{\theta}(\mathbf{x}_{t-1}^{i-1},t-1)$
@@
12:  for $t=0,\cdots,1$ ($T$ steps) do
13:     $\mathbf{\hat{x}}_{t+1}=\mathbf{x}_{t}+(\sigma_{t+1}-\sigma_{t})v_{\theta}(\mathbf{x}_{t},t)$
14:     $\epsilon_{t}=\mathbf{x}_{t+1}-\mathbf{\hat{x}}_{t+1}$
15:  end for
16:  Output: The inversion trajectory $\{\mathbf{x}_{t}\}$ and $\{\epsilon_{t}\}$
@@ 5 Flexible Invariance Control with AdaLN
Module design. In the investigation of MM-DiT, our key observation is that the text features within the adaptation layer normalization (AdaLN) closely correspond to the image semantics such as pose, quantity, and object types. Thus, we can inject unchanged text features from the original text into the target and keep the features of the edited text intact. This will only change regions indicated by text changes. Concretely, as depicted in Figure 2, we denote the source and target text features of the AdaLN before the attention in the MM-DiT block as $M^{s},M^{t}\in\mathbb{R}^{j\times d}$ where $j$ represent the number of tokens of the text prompt and $d$ represent the text feature dimension. We define the token-aware Map function $\mathbf{Map}(M^{s},M^{t},\mathcal{P}_{s},\mathcal{P}_{t})$ to replace the features of unedited tokens in the target prompt $\mathcal{P}_{t}$ with corresponding tokens in the source prompt $\mathcal{P}_{s}$. Such operation is conducted starting from the beginning to the timestep $S$.
$\mathbf{Map}(M^{s},M^{t},\mathcal{P}_{s},\mathcal{P}_{t}):=\left\{\begin{aligned} \hat{M}^{t}&,if\ t<S,\\ M^{t}&,otherwise.\end{aligned}\right.$ (12)
The proposed module connects the features to the changes in the text prompt and can flexibly control the invariance without hindering the editing. In practice, we build such an editing module based on the MM-DiT Stable Diffusion 3.5. We show the potential of editing in various scenarios in Figure 4. The results show that the replaced text features $\hat{M}^{t}$ in AdaLN adapt to different editing scenarios. To complete our argument, we also present the investigation of Self-Attention of MM-DiT for invariance control in Figure 5. The observations are similar to previous experiences in U-Net. Attention injection in too many timesteps may prevent some non-rigid editing effects. In contrast, we keep the injection timestep $S$ the same for all editing types, and such a unified and stable timestep does not hinder non-rigid editing. Please refer to Appendix for details.
Figure 5: Invariance control of self-attention with different time periods. We inject different components (Q,K,V) of self-attention with different time periods to verify the ability to preserve the unedited contents (Appendix). Here we show injecting V value more than 20% time steps can hinder the non-rigid editing ‘sitting to standing’.
@@ 6.1 Setup
Baselines and implementations. We mainly compare our methods with previous state-of-the-art turning-free image editing methods: P2P [10], PnP [35], MassCtrl [4], InfEdit [39]. All these methods use attention maps in U-Net to control the invariance for editing. Specifically, MasaCtrl and InfEdit address the issue of non-rigid editing. Additionally, we also include the instruction-based editing method InsP2P [3] which is retrained with the triplet of before-and-after image pairs and the editing instruction. For all methods, we use their official implementations for evaluation. We implement our method based on the Stable Diffusion 3.5 and use the basic Euler sampler of rectified flow for sampling and inversion. We use 30 steps for inversion and choose the CFG as 1 and 2 for inversion and editing respectively. We set the fixed-point iteration number as 3. For methods based on diffusion U-Net and DDIM sampling, we use 50 as inversion steps and choose CFG as 7, which follows their default settings.

### v4

@@ Abstract
Leveraging the large generative prior of the flow transformer for tuning-free image editing requires authentic inversion to project the image into the model’s domain and a flexible invariance control mechanism to preserve non-target contents. However, the prevailing diffusion inversion performs deficiently in flow-based models, and the invariance control cannot reconcile diverse rigid and non-rigid editing tasks. To address these, we systematically analyze the inversion and invariance control based on the flow transformer. Specifically, we unveil that the Euler inversion shares a similar structure to DDIM yet is more susceptible to the approximation error. Thus, we propose a two-stage inversion to first refine the velocity estimation and then compensate for the leftover error, which pivots closely to the model prior and benefits editing. Meanwhile, we propose the invariance control that manipulates the text features within the adaptive layer normalization, connecting the changes in the text prompt to image semantics. This mechanism can simultaneously preserve the non-target contents while allowing rigid and non-rigid manipulation, enabling a wide range of editing types such as visual text, quantity, facial expression, etc. Experiments on versatile scenarios validate that our framework achieves flexible and accurate editing, unlocking the potential of the flow transformer for versatile image editing. Project Page is here.
@@ 1 Introduction
To address the aforementioned issues, we systematically investigate flow inversion with the Euler method and invariance modeling in MM-DiT. For inversion, we reveal that the reverse ordinary differential equation (ODE) with the Euler method has a similar structure to DDIM inversion but is more affected by the approximation error between two consecutive states. This error makes Euler inversion diverge more from the original image than DDIM inversion in diffusion and significantly deteriorates the editing ability even if the image can be recovered by optimization. So, we proposed a two-stage flow inversion. First, we use the fixed-iteration technique to reduce the approximation error when estimating the flow velocity and get a basic inversion with the Euler method. This makes inversion approximates the authentic generation process. Then based on such inversion, we only need to add mild compensation at each denoising step to recover the exact original image. Since the compensation at each denoising step is small, the inversion trajectory is close to the model’s original domain rather than overfitting one image, thereby preserving the editing ability.
@@ 2 Related Work
Image editing via diffusion models. Earlier editing methods fine-tune the model with the original image and the text prompt [18, 55]. Later approaches mainly fall into three categories. Truncation methods analyze latent noise predictions of the model and filter out undesired editing parts to preserve the non-target regions [2, 19]. The instruction-training methods [4, 53, 8, 39, 26, 10] collects a triplet of original and edited images and the editing instruction to re-train the diffusion model to be an editing model, which can be difficult to scale up. The inversion-based methods [50, 11, 30, 43, 5] are tuning-free and first convert the original image into the initial noisy latent and regenerate the edited image with the edited prompt while preserving the invariance (i.e., unchanged content from the original image). Others further improve the applicable ability by combining the large language model [15, 20] and enrich the editing forms [40, 51, 21, 24].  Hu et al. [14] explores the FMs but based on U-ViT. [37, 46] explore the DiT yet they show limited tradeoffs in preserving invariance and editing in versatile non-rigid editing types.
@@
Inversion of diffusion for real image editing. The inversion that converts the image into the initial noisy latent is pivotal in tuning-free editing. One branch leverages DDIM [41] to invert the ODE process. However, the classifier-free guidance (CFG) in editing deteriorates the accumulated error in reversing ODE. Various mitigation solutions explore optimizing the null-text embedding [28], adding corrections [17], better estimation of latents [7], and negative prompt [27]. Another branch adopts the stochastic differential equation (SDE) instead of ODE to mitigate the error by manipulating random noise [3, 16]. Besides these heuristic solutions, others explore the mathematical exact inversion by reformulating the sampling process [44, 13, 45]. Besides these nonparametric solutions, TurboEdit [47] models the inversion process with a neural network to reduce the inversion time. Unlike these diffusion-oriented inversions, we specifically unveil the deficiency of flow inversion and study efficient inversion based on the velocity field of flow models.
@@
Invariance control in diffusion-based image editing. The invariance control preserves the original image’s non-target contents during editing. The instruction-based methods explicitly add the original image features as the condition and retrain the T2I model into a text-guided image-to-image (TI2I) model [4, 8, 54, 52]. In the tuning-free paradigm, attention-based methods [11, 43, 46] inject attention maps from the original image to preserve the original contents but may struggle with non-rigid editing such as layout and pose. Later methods [5, 49, 46] inject K and V values of attentions in the later diffusion process to adapt to the layout change but this deteriorates the fidelity for large changes. The refinement approach [2, 19] filters out components of the predicted noise corresponding to the non-target regions.  Rout et al. [37] regulates the flow inversion with the original image. However, these approaches require careful tuning of hyperparameters for different editing types, and most are based on diffusion and U-Net. In contrast, we systematically investigate inversion for rectified flow and flexible invariance control based on MM-DiT to support rigid and non-rigid editing, enabling versatile editing types.
@@
Figure 2: Framework of rectified flow inversion and editing. Left: The two-stage inversion for rectified flow. A basic inversion trajectory is first constructed to resemble the generation process, and then mild compensation is added to the velocity to recover the image exactly. Right: Invariance control with text feature replacement within AdaLN during sampling. In the target branch, the unchanged text features are replaced with features from the original image while the edited text features remain intact.
@@ 3 Preliminary
We aim to design a tuning-free framework based on the flow transformer for versatile editing. In Section 3, we briefly introduce the rectified flow and DDIM. Then, in Section 4, we analyze the deficiency of Euler inversion and propose the two-stage flow inversion. In Section 5, we present the image invariance investigation in MM-DiT centered on AdaLN. Last, we summarize the full algorithm.
@@
DDIM. DDIM [41] is the implicit probabilistic model trained with DDPM [12] objective and has the formulation below where $\mathbf{x}_{t}$ is also the noisy data sample, $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$ is the noise prediction, $\mathbf{f}_{\theta}(\mathbf{x}_{t},t)$ is the predicted original data sample.
@@ 4.1 Deficiency and Relation of Euler to DDIM
Essentially, if we rescale $\mathbf{x}_{t}$ with $\frac{\mathbf{x}_{t}}{\sqrt{\alpha_{t}}}$ and rewrite Eq. 5, DDIM can also be treated as the first-order ODE solver as Euler above. The corresponding inversion is Eq. 9.
$\frac{\mathbf{x}_{t}}{\sqrt{\alpha_{t}}}$ $\approx\frac{\mathbf{x}_{t+1}}{\sqrt{\alpha_{t+1}}}+(\sqrt{\frac{1-\alpha_{t}}{\alpha_{t}}}-\sqrt{\frac{1-\alpha_{t+1}}{\alpha_{t+1}}})\mathbf{s}_{\theta}(\mathbf{x}_{t+1},t)$ (9)
However, despite the similar formulation of Euler and DDIM inversion, the performance of Euler inversion is significantly worse than DDIM inversion, even if the Euler inversion is evaluated on the more powerful Stable Diffusion 3.5 vs SD 1.5 as shown in Figure 3. Such phenomena are not expected and also make optimization methods [28, 17, 3] in diffusion invalid for flow inversion.
@@
We argue this is mainly due to that the Euler inversion for the rectified flow transformer suffers more from the approximation error. To reduce it, we need to mitigate the gap between $v_{\theta}(\mathbf{x}_{t+1},t)$ and $v_{\theta}(\mathbf{x}_{t},t)$. Idealy, we expect to use the exact inversion in Eq. 7 to get the initial noisy latent $\mathbf{x}_{0}$ and the transition trajectory $\{\mathbf{x}_{1},...,\mathbf{x}_{t}\}$. Such an inversion trajectory is close to the trajectory in the generating process and requires minimal optimization. Thus, it fits the prior distribution of the model and has good editing ability.
@@ 4.2 Two-stage Flow Inversion
Our principle is to find an initial latent and inversion trajectory that is as close to the authentic generation process as possible, trying to avoid only overfitting the original image. This has two-fold benefits: 1. Editing friendly. The latent and trajectory fit into the prior distribution of the model and text-to-image alignment is authentic so that the image can be easily manipulated by changing the text prompt. 2. Easy to preserve invariance and fidelity. If the inversion trajectory deviates from the model prior, it requires more effort, such as self-attention injection to preserve the invariance and fidelity. However, injecting too much self-attention from the original image can hinder the editing. Thus, we propose a two-stage flow inversion. First, we construct a pivotal inversion close to the generating process. Second, we manually add compensation at every timestep to eliminate the left mild error.
@@
Besides, ideally, the velocity $v_{\theta}$ is $\mathbf{x}_{1}-\mathbf{x}_{0}$ according to Eq. 3 that is a stable value. This also strengthens the motivation of fixed-point iteration, and we can use this property to average the velocities to get a more stable velocity prediction. Practically, this is also equivalent to averaging the latent series $\{\mathbf{x}_{t}^{i}\}_{i=1}^{I}$. We compare the latents with and without averaging at different steps in Figure 4 and show that averaged latents are closer to the original image since the fixed-point iteration may not converge monotonically and wander around the true fixed point.
Figure 4: Visualization of latents $\mathbf{x}_{t}$ w / w.o averaging.
@@
1:  Input: origin image $\mathbf{x}_{1}$, iteration steps $I$, inversion steps $T$, velocity model $v_{\theta}$, before after prompts $\mathcal{P}_{b}$, $\mathcal{P}_{a}$, timestep $S$.
@@
6:        $\mathbf{x}_{t-1}^{i}=\mathbf{x}_{t}+(\sigma_{t-1}-\sigma_{t})v_{\theta}(\mathbf{x}_{t-1}^{i-1},t-1,\mathcal{P}_{s})$
@@
12:  Initialize before $\mathbf{x}^{b}_{0}$ and after $\mathbf{x}^{a}_{0}$ with $\mathbf{x}_{0}$.
13:  for $t=0,\cdots,1$ ($T$ steps) do
14:     if $t<S$ then $\mathbf{Map}(M^{b},M^{a},\mathcal{P}_{b},\mathcal{P}_{a})$
15:     $\mathbf{\hat{x}}^{b}_{t+1}=\mathbf{x}^{b}_{t}+(\sigma_{t+1}-\sigma_{t})v_{\theta}(\mathbf{x}^{b}_{t},t,\mathcal{P}_{b})$
16:     $\epsilon_{t}=\mathbf{x}^{b}_{t+1}-\mathbf{\hat{x}}^{b}_{t+1}$
17:     $\mathbf{x}^{a}_{t+1}=\mathbf{x}^{a}_{t}+(\sigma_{t+1}-\sigma_{t})v_{\theta}(\mathbf{x}^{a}_{t},t,\mathcal{P}_{a})+\epsilon_{t}$
18:  end for
19:  Output: The edited after image $\mathbf{x}^{a}_{1}$
@@ 5 Flexible Invariance Control with AdaLN
Module design. In the investigation of MM-DiT, our key observation is that the text features within the adaptation layer normalization (AdaLN) closely correspond to the image semantics such as pose, quantity, and object types. Thus, we can inject unchanged text features from the original text into the target and keep the features of the edited text intact. This will only change regions indicated by text changes. Concretely, as depicted in Figure 2, we denote the before and after text features of the AdaLN before the attention in the MM-DiT block as $M^{b},M^{a}\in\mathbb{R}^{j\times d}$ where $j$ represent the number of tokens of the text prompt and $d$ represent the text feature dimension. We define the token-aware Map function $\mathbf{Map}(M^{b},M^{a},\mathcal{P}_{b},\mathcal{P}_{a})$ to replace the features of unedited tokens in the after prompt $\mathcal{P}_{a}$ with corresponding tokens in the before prompt $\mathcal{P}_{b}$. Such operation is conducted starting from the beginning to the timestep $S$.
$\mathbf{Map}(M^{b},M^{a},\mathcal{P}_{b},\mathcal{P}_{a}):=\left\{\begin{aligned} \hat{M}^{a}&,if\ t<S,\\ M^{a}&,otherwise.\end{aligned}\right.$ (13)
The proposed module connects the features to the changes in the text prompt and can flexibly control the invariance without hindering the editing. In practice, we build such an editing module based on the MM-DiT Stable Diffusion 3.5. We show the potential of editing in various scenarios in Figure 5. The results show that the replaced text features $\hat{M}^{a}$ in AdaLN adapt to different editing scenarios. To complete our argument, we also present the investigation of Self-Attention of MM-DiT for invariance control in Figure 6. The observations are similar to previous experiences in U-Net. Attention injection in too many timesteps may prevent some non-rigid editing effects. In contrast, we keep the injection timestep $S$ the same for all editing types, and such a unified and stable timestep does not hinder non-rigid editing. Please refer to Appendix Sec: 11 for details.
Figure 6: Invariance control of self-attention with different time periods. We inject different components (Q,K,V) of self-attention with different time periods to verify the ability to preserve the unedited contents (Appendix Sec: 11). Here we show injecting V value more than 20% time steps can hinder the non-rigid editing ‘sitting to standing’.
@@ 6.1 Setup
Baselines and implementations. We mainly compare our methods with previous state-of-the-art turning-free image editing methods: P2P [11], PnP [43], MassCtrl [5], InfEdit [49], RFinv [37]. The first four use attention maps in U-Net to control the invariance for editing. Specifically, MasaCtrl and InfEdit address the issue of non-rigid editing. RFinv [37] is close to our work exploring flow-based models for tuning-free editing, but is based on Flux-dev. Additionally, we also include the instruction-based editing method InsP2P [4], which is retrained with the triplet of before-and-after image pairs and the editing instruction. We follow their official implementations for evaluation for all methods. We implement our method based on the Stable Diffusion 3.5 and use the basic Euler sampler of rectified flow for sampling and inversion. We use 30 steps for inversion and choose the CFG as 1 and 2 for inversion and editing respectively. We set the fixed-point iteration number as 3. For methods based on diffusion U-Net and DDIM sampling, we use 50 as inversion steps and choose CFG as 7, which follows their default settings.